4g H 2 ,32g O 2 , 14g N 2 and 11g CO 2 are taken in a bulb of 500ml. Which one of these has maximum active mass? (1)N 2 (2)O 2 (3)H 2 (4)CO 2

4g H2,32g O, 14g Nand 11g CO2 are taken in a bulb of 500ml. Which one of these has maximum active mass?
(1)N2
(2)O2
(3)H2
(4)CO2

Grade:11

1 Answers

Vikas TU
5 years ago
moles of H2 = 4/2 = 2
moles of N2 = 14/14 = 1
moles of O2 = 32/32 = 1
moles of CO2 = 11/44 = ¼ = 0.25
 
Hence acyive mass is proportionl to the number of moles therfore, H2 has maximum number of active mass.